the equation is 2x+y = 12. the question is write an equation of a line passing through the point 2,4. and that is perpencidular to the given equation. I had come up with y = 1/2 + 5. Using the inverse of the original slope which was
-2. its marked wrong, I dont know why

Answer by Alan3354(69443)

the equation is 2x+y = 12. the question is write an equation of a line passing through the point 2,4. and that is perpencidular to the given equation. I had come up with y = 1/2 + 5. Using the inverse of the original slope which was
-2.
===============
2x+y = 12 has a slope of -2, that's correct.
Perpendicular lines has a slope of 1/2
-----
Use y = mx + b and the point to find b
4 = (1/2)*2 + b
b = 3
-----
--> y = (1/2)x + 3
-----------------------
Your equation had no x, too.